### v1

# Five parameters are all you need

Abstract
The standard cosmological model, with its six independent parameters, has proven remarkably successful in describing the evolution of the Universe. One of these parameters, the optical depth to reionization $\tau_{\mathrm{reio}}$, represents the scatterings that Cosmic Microwave Background (CMB) photons will experience after decoupling from the primordial plasma as the intergalactic medium transitions from neutral to ionized. $\tau_{\mathrm{reio}}$ depends on the neutral hydrogen fraction $x_{\mathrm{HI}}$, which, in turn, should theoretically depend on cosmology. We present a novel method to establish the missing link between cosmology and reionization timeline. We discover the timeline has a universal shape well described by the Gompertz mortality law, applicable to any cosmology within our simulated data. This enables us to map cosmology to reionization using symbolic regression and to treat $\tau_{\mathrm{reio}}$ as a derived parameter. Reanalyzing CMB data with our universal $x_{\mathrm{HI}}$ tightens the constraint on $\tau_{\mathrm{reio}}$ by more than one order of magnitude to $\approx 1\%$ and reduced the error on the amplitude of the primordial fluctuations by a factor of 2.5 compared to Planck’s PR3 constraint. While our results rely on the astrophysical assumptions in our simulations, the methodology itself is independent of these assumptions; after all, $\tau_{\mathrm{reio}}$ is fundamentally a function of cosmology.
The $\Lambda$CDM cosmological model has proven extremely effective in predicting the evolution of our Universe, relying on only six parameters 1. In particular, it explains the transition from a predominantly neutral state in the early stages to the familiar ionized intergalactic medium (IGM) observed in our relatively nearby surroundings. This transition is known as cosmic reionization. Despite a comprehensive understanding of the astrophysical principles governing this transition, uncertainties persist regarding its precise timeline 2. The advent of the James Webb Space Telescope (JWST) 3 represents a pivotal moment, substantially bolstering our ability to directly constrain the evolution of the neutral hydrogen fraction $x_{\mathrm{HI}}$. This progress is being driven by the JWST’s enhanced detection capabilities, enabling the observation of high-redshift quasars 4 and high-redshift galaxies 5, 6, 7, 8.
Reionization leads to scattering of Cosmic Microwave Background (CMB) photons by free electrons, disrupting the CMB angular power spectra ($C_{\ell}$). This scattering suppresses the signal at scales smaller than the Hubble scale at reionization (approximately $\ell>10$) 9 due to the optical depth $\tau_{\mathrm{reio}}$. Additionally, it introduces a new signal in the polarization of CMB photons at large angular scales 1, that is $\propto\tau_{\mathrm{reio}}$ in $C^{TE}_{\ell}$, the cross-correlation of the $E$-mode polarization with the temperature (intensity), and is $\propto\tau_{\mathrm{reio}}^{2}$ in $C^{EE}_{\ell}$, the $E$-mode polarization angular auto power spectrum. Consequently, heightened sensitivity to CMB polarization becomes crucial for mitigating the degeneracy between $\tau_{\mathrm{reio}}$ and other cosmological parameters, particularly $A_{\mathrm{s}}$, the amplitude of the primordial scalar power spectrum, and $r$, the ratio of tensor-to-scalar modes 10.
$-\ell$ polarization data is crucial to determine $\tau_{\mathrm{reio}}$; however, the measurement of such a weak signal ($\sim 10^{-2}\mu$K2) demands superb systematic and foreground control9. Furthermore, anomalous measurements in $C^{TE}_{\ell}$ at low multipoles1 could indicate concerns to the cosmological interpretations at these angular scales. Ultimately, this challenging measurement may require adopting a comprehensive Bayesian framework to jointly consider cosmology, astrophysics, and instrument systematics11. Figure 1 illustrates current representative constraints on $\tau_{\mathrm{reio}}$.
[…]
Given the challenges posed by $\tau_{\mathrm{reio}}$ in CMB analyses and the anticipated advancements in constraining the reionization timeline 15, 16, now is an opportune moment to reassess its role. In theory, cosmic reionization is uniquely determined given a specific cosmology, i.e. by the other five cosmological parameters. Specifically, the evolution of the global fraction of neutral hydrogen can be written as
$x_{\mathrm{HI}}(a)=f(a;\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}})\quad\Rightarrow\quad\tau_{\mathrm{reio}}=g(\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}}),$ (1)
where $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, and $\Omega_{\mathrm{m}}$ are the tilt of the primordial power spectrum, dimensionless Hubble constant, and present baryon and matter density fractions, respectively. $\sigma_{8}$ is the present linear rms relative density fluctuation in a sphere of radius $8\,h^{-1}$pc. This parametrization chosen for convenience fully determines $x_{\mathrm{HI}}$, but the incomplete understanding of cosmic reionization obscures this mapping, necessitating the introduction of $\tau_{\mathrm{reio}}$ in CMB analyses. However, our understanding of the astrophysical processes governing reionization has significantly improved 17, 18, 19, 20 since the inclusion of $\tau_{\mathrm{reio}}$ became a standard practice. Ongoing and forthcoming observations promise to further our understanding and reduce inherent modeling uncertainties. Motivated by these developments, we use symbolic regression (SR) 21 to construct a mapping between cosmology and reionization timeline, aiming to demote $\tau_{\mathrm{reio}}$ from an independent to a derived cosmological parameter, and simultaneously tightening constraints in cosmological parameters.
Eq. 1 introduces a novel avenue to constrain cosmology by examining the dependence of $x_{\mathrm{HI}}(z)$ on cosmological parameters. This mapping can enhance parameter constraints and shed light on reionization astrophysics. It also aids ongoing efforts in parametrizing cosmic reionization models 22, 23, 12 by including the cosmological dependence of $x_{\mathrm{HI}}$.
Here, we present a universality in the neutral hydrogen time evolution, and derive through SR its dependence on cosmology on simulated reionization histories from 21cmFAST 19. We integrate this shape into CLASS 24, a popular Boltzmann solver for CMB analyses. We then evaluate the modified CLASS alongside Cobaya 25, a speed-aware sampler26, 27111With adaptive covariance learning and fast-dragging as in 28, enabling larger steps in slow parameters via intermediate transitions of fast parameters., showcasing its ability to recover parameter constraints from CMB data, including ‘TTTEEE’ + lensing likelihoods 29, 30 (see Figure 9). Finally, we demonstrate cosmological gains by computing $\tau_{\mathrm{reio}}$ as a derived parameter using our approach – Eq. 1 – compared to sampling over $\tau_{\mathrm{reio}}$ using the conventional $\tanh$ model 31. We summarize our strategy in Figure 2.
Figure 2: Strategy to demote $\tau_{\mathrm{reio}}$ to derived parameter. a) Sobol sampling of $\bm{\theta}$ comprising 5 cosmological and 1 astrophysical parameters (see Figure 5 in Extended Data). b) Simulated $x_{\mathrm{HI}}$ timelines as a function of $\bm{\theta}$ and scale factor $a$. c) With symbolic regression, we optimize the mapping from $\bm{\theta}$ to the rescaling parameters that bring the universality. d) We model the universal shape (upper panel) as a composition of the Gompertz function and a low-degree polynomial (lower panel). e) Planck CMB data we re-analyze. f) We infer the parameter constraints using Monte Carlo Markov Chain (MCMC).
To achieve our goal outlined in Eq. 1, we construct a universal shape for $x_{\mathrm{HI}}$ using 128 Sobol samples of 21cmFAST simulations (see Simulations and Figure 5 in the Extended Data). All $x_{\mathrm{HI}}(a)$ profiles share this shape, with differences between scenarios being mere translations and rescalings. Reionization causes $x_{\mathrm{HI}}$ to reduce from near 1 to effectively 0 via a sigmoid transition. The standard $\tanh$ function is symmetric in nature and not flexible enough to provide the early start and rapid completion suggested by reionization simulations 22, 32. The Gompertz curve, an asymmetric sigmoid function often used to analyze age-dependent human mortality 33, proves a good model for the survival of neutral hydrogen too. Its expected accelerated increase in mortality with age resembles the expectation for the percolation of ionized HII bubbles during the end stages of reionization.
One way to uncover the universality is to view each $x_{\mathrm{HI}}$ scenario as a cumulative probability distribution (CDF) in $-\ln a$. With this insight, we can translate and rescale each timeline using the mean and variance of its corresponding probability density function (PDF), i.e. $-\mathrm{d}x_{\mathrm{HI}}/\mathrm{d}\ln a$, and discover the existence of a universal shape followed by all scenarios. Therefore, cosmology only impacts the translation and rescaling parameters of each timeline, not its shape.
However, with the PDF trick, some $x_{\mathrm{HI}}$’s can deviate artificially from universality, due to their incomplete reionization given our broad range of simulated scenarios. To address this, we adopt a better approach to jointly fit the global shape and the 2 individual parameters of each $x_{\mathrm{HI}}$. Our shape model constitutes the Gompertz function composed with a 5th-degree polynomial, in the translated-and-rescaled time $\ln\tilde{a}\equiv\beta(\ln a-\ln\alpha)$. And we are free to set the polynomial constant to 0 and its linear coefficient to 1 by utilizing their respective degeneracy with $\ln\alpha$ and $\beta$.
The complete model parametrizes the HI evolution as follows (also see Figure 6 and Shape universality and modeling in the Extended Data):
$x_{\mathrm{HI}}(\tilde{a})$ $=\mathrm{gomp}\bigl{(}P_{5}(\tilde{a})\bigr{)}\equiv\exp\bigl{[}-\exp\bigl{(}P_{5}(\tilde{a})\bigr{)}\bigr{]},$ (2)
$P_{5}(\tilde{a})$ $={\textstyle\sum}_{m=0}^{5}\,c_{m}\ln^{m}\!\tilde{a},$ (3)
$\bm{c}$ =$=\{0,1,0.1503,0.04850,0.005261,0.0002182\},$
$\tilde{a}(a;\bm{\theta})$ $=\Bigl{[}\frac{a}{\alpha(\bm{\theta})}\Bigr{]}^{\beta},$ (4)
where $\bm{\theta}$ denotes 6 astrophysical and cosmological parameters, $\alpha(\bm{\theta})$ is the power-law pivot (or logarithmic translation), and $\beta=8.290$ is the rescaling tilt, which, according to our 21cmFAST simulations, appears constant. This lack of cosmological dependence may stem from 21cmFAST’s treatment of reionization astrophysics. See Helium reionization for specifics on implementing HeI and HeII reionization.
Before fully leveraging our formalism to extract the cosmological dependence in the rescaling of Eq. 2 and relaxing the need for $\tau_{\mathrm{reio}}$ in CMB analyses, we first implement the Gompertz shape with independent $\tau_{\mathrm{reio}}$ in CLASS and confirm its agreement with the conventional $\tanh$ model (gomp 0 and $\tanh$ 0 in Table 1). Using Planck PR3 likelihoods ‘TTTEEE’ 29 and CMB lensing 30, we sample typical cosmological parameters with Cobaya 25, including $\tau_{\mathrm{reio}}$. Given a proposal for $\tau_{\mathrm{reio}}$, we determine the corresponding reionization timeline using bisection by varying $\ln\alpha$ for gomp 0, while for $\tanh$ 0, the reionization midpoint $z_{\mathrm{re}}$ is the tuning parameter. The sampler runs until the Gelman-Rubin statistic 34 satisfies $R-1<0.2$ for the between-chain variance of the confidence intervals. We repeat this for $\tanh$ 0 and verify the agreement between the two models.
[…]
Gompertz-polynomial-shaped reionization can reproduce standard CMB analyses. Now, we move to establish the connection between the universal shape for $x_{\mathrm{HI}}$ and the rescaling of a given reionization scenario. We refer to this model as gomp 1. This rescaling is naturally a function of cosmology. For example, a larger density of matter $\Omega_{\mathrm{m}}$ results in deeper potential wells, accelerating structure formation and increasing the number of ultraviolet photons driving the reionization process. We employ PySR, an SR package, to establish the cosmological dependence of the rescaling in Eq. 4.
While PySR initially guided us towards the Gompertz curve when directly regressing $x_{\mathrm{HI}}$, the final analysis only uses it to regress the pivot and tilt instead. We fit their values jointly with the polynomial coefficients as described above, and feed them as labels to the genetic algorithm to find the best analytic expression (see Symbolic regression in Extended Data for our definition of best). Using PySR we derived the following pivot as a function of cosmology
$\ln\alpha(\bm{\theta})=(h^{\Omega_{\mathrm{m}}}+\sigma_{8})(\Omega_{\mathrm{b}}-n_{\mathrm{s}}-\Omega_{\mathrm{m}}),$ (5)
which, interestingly, is independent of the astrophysical parameter $\zeta_{\mathrm{UV}}$ (see Simulations in Extended data). This ionization efficiency modulates the timing of reionization in our simulations by regulating the abundance of photons that escape into the IGM. Its absence in Eq. 5 suggests that it either plays a minor role or its contribution can be mimicked by the cosmological parameters.
Eqs. 4 and 5 imply that higher values of $n_{\mathrm{s}}$ hasten reionization by boosting power on small scales, fostering a greater abundance of ionizing sources and earlier completion 15. Note that our 21cmFAST simulations assume that faint galaxies are the primary drivers of reionization. Similarly, larger $\Omega_{\mathrm{m}}$ and $\sigma_{8}$ primarily expedite reionization by enhancing structure formation. Surprisingly, Eq. 5 suggests that higher $\Omega_{\mathrm{b}}$ delays reionization, likely due to the increased abundance of HI in the intergalactic medium requiring more ionizing photons.
We note that within the prior range of our 21cmFAST simulations (see Simulations) and their corresponding astrophysics of reionization, the mapping derived from SR is not unique. Additional details and results using an alternative mapping – gomp 2 – are presented in Alternative mapping of the Extended Data. Nonetheless, our results are robust and independent of the choice of mapping.
We implement Eq. 5 in our Gompertz CLASS, which given the cosmological parameters determines the pivot value and consequently the reionization history, $\tau_{\mathrm{reio}}$, and corresponding CMB angular power spectra. This gomp 1 model eliminates the need to sample over $\tau_{\mathrm{reio}}$ (or $z_{\mathrm{re}}$), requiring only five cosmological parameters, e.g. $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}}\}$. We use Cobaya to re-analyze the same CMB data, showcasing the full strength of our approach.
Figure 3: Analysis of CMB data with reionization as a function of cosmology. The green contours represent our results using the Gompertz reionization model with Eq. 5, eliminating the need to sample over any reionization parameter. The blue contours correspond to the results obtained using the conventional $\tanh$ model, while the relevant Planck constraints 1 are depicted with gray lines for reference.
Figure 3 underscores the impact of our universally-shaped Gompertz reionization model. This approach tightens the constraint on the optical depth to $\approx 1\%$, a remarkable improvement compared to $>10\%$ with the $\tanh$ prescription. Furthermore, the constraint on $A_{\mathrm{s}}$ improves dramatically since the TT data is no longer significantly hampered by the degeneracy between $A_{\mathrm{s}}$ and $\tau_{\mathrm{reio}}$. The error on $A_{\mathrm{s}}$ decreases by an impressive factor of 2.5 compared to Planck’s results 1. Overall, we recover tighter constraints across the board compared to Planck. See Figures 8 and 1 in the Extended Data for details.
[…]
Our findings for the timeline of reionization align with late reionization scenarios, which are supported by high-$z$ Lyman-$\alpha$ observations 35, 49. However, recent discoveries by JWST indicate the presence of massive, bright galaxies at early redshifts $z\sim 10$ 5, 6, 7. The presence of these early galaxies suggests a potential preference for brighter galaxies to drive reionization, a role that in our 21cmFAST simulations was attributed to a population of faint galaxies.
Furthermore, our results are influenced by the semi-numerical prescription employed by 21cmFAST to ionize the IGM, which, while efficient and swift, could bias our findings. Moreover, our exploration within the astrophysical framework of 21cmFAST has been limited to varying the ionization efficiency (see Simulations in the Extended Data). Therefore, a more comprehensive exploration is warranted to ensure, for instance, that the rescaling tilt $\beta$ remains a constant. Additionally, a valuable exercise to refine the inherent relationship between cosmological parameters and reionization history would involve using more realistic, albeit slower, reionization models. One such option is to use the THESAN simulations 18, which are hydrodynamical simulations incorporating radiative transfer.
Methodology
Simulations
To establish the universality of our proposed Gompertz model for the neutral hydrogen reionization and its relationship with the cosmological parameters, we conducted 128 21cmFASTv3 simulations to generate the corresponding $x_{\mathrm{HI}}(z;\bm{\theta})$ profiles. Our parameter space include $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},h,\Omega_{\mathrm{b}},\Omega_{\mathrm{m}},\zeta_{\mathrm{UV}}\}$, comprising five cosmological and one astrophysical parameters. The selection of $\sigma_{8}$ instead of $A_{\mathrm{s}}$ is imposed by the input requirements of 21cmFAST [^2] Alternatively, one could use 21cmFirstCLASS50 to avoid $\sigma_{8}$.. We use a scrambled Sobol sequence 51, 52 to sample quasi-uniformly in the $\bm{\theta}$-space, within the following ranges:
$\sigma_{8}$ $\in(0.74,0.90),$ $\quad n_{\mathrm{s}}$ $\in(0.92,1.00),$ $\quad h$ $\in(0.61,0.73),$
$\Omega_{\mathrm{b}}$ $\in(0.04,0.06),$ $\quad\Omega_{\mathrm{m}}$ $\in(0.24,0.40),$ $\quad\zeta_{\mathrm{UV}}$ $\in(15,30).$ (6)
Their 1D and 2D projections in Figure 5 illustrate the sample uniformity in parameter space. Our 21cmFAST simulations have a 300 comoving Mpc box size and $768^{3}$ ($256^{3}$) cells for the matter (HI) field. We maintain most options in their default values and extract the neutral hydrogen fraction using lightcone.global_xH.
Figure 5: Sobol sampling of our $x_{\mathrm{HI}}$ profiles, in 6D parameter space of $\sigma_{8}$, $n_{\mathrm{s}}$, $h$, $\Omega_{\mathrm{b}}$, $\Omega_{\mathrm{m}}$, and $\zeta_{\mathrm{UV}}$, Each point in the lower triangular panels corresponds to the 2D projection of a 21cmFAST run, while the diagonal panels show the 1D cumulative histograms for each parameter. These 1D and 2D projections demonstrate the uniformity of our sampling of the parameter space within the prior range of Simulations.
The ionization efficiency $\zeta_{\mathrm{UV}}$ governs the ability of ultraviolet photons to escape their parent galaxies and ionize the IGM. An increase in $\zeta_{\mathrm{UV}}$ leads to an earlier completion of HI reionization. Due to significant uncertainties surrounding $\zeta_{\mathrm{UV}}$, we opted to use a constant value in each simulation. Note that this choice was made for simplicity, and a more realistic assumption could be that $\zeta_{\mathrm{UV}}$ is a function of halo mass 53 or redshift. However, our findings suggest that $\zeta_{\mathrm{UV}}$ plays a minor role compared to the cosmological parameters. In fact, the best performing symbolic regression expressions (see Symbolic regression for the selection of best models) do not use $\zeta_{\mathrm{UV}}$.
Helium reionization
The early intergalactic medium is primarily composed by neutral hydrogen and helium. Neutral helium (HeI) loses is first electron at the same time as neutral hydrogen (HI) gets ionized 54. However, there is a second reionization that occurs around $z\sim 3$ where Helium (HeII) loses its remaining electron.
CMB photons will scatter off any free electrons, therefore both helium reionizations contribute to the Thomson optical depth to reionization $\tau_{\mathrm{reio}}$, although HeII ionization contribute relatively little in comparison to HI and HeI ionizations 55.
To include the impact of the first helium reionization in our Gompertz CLASS, we assume it follows that of HI as done in the $\tanh$ model, i.e. the free electron fraction $x_{\mathrm{e}}$ is given by
$x_{\mathrm{e}}$ $=\Bigl{(}1+\frac{n_{\mathrm{He}}}{n_{\mathrm{H}}}-x^{\mathrm{rec}}_{\mathrm{e}}\Bigr{)}x_{\mathrm{e}}^{\mathrm{gomp}}+x^{\mathrm{rec}}_{\mathrm{e}}$
$=\Bigl{(}1+\frac{Y_{\mathrm{He}}}{C(1-Y_{\mathrm{He}})}-x^{\mathrm{rec}}_{\mathrm{e}}\Bigr{)}x_{\mathrm{e}}^{\mathrm{gomp}}+x^{\mathrm{rec}}_{\mathrm{e}},$ (7)
where $n_{\mathrm{He}}/n_{\mathrm{H}}$ is the helium to hydrogen number density ratio, $C\equiv m_{\mathrm{He}}/m_{\mathrm{H}}\approx 4$ is their mass ratio, $Y_{\mathrm{He}}$ is the helium mass fraction, $x_{\mathrm{e}}^{\mathrm{gomp}}$ corresponds to the contribution of free electrons due to Eq. 2, and $x^{\mathrm{rec}}_{\mathrm{e}}\approx 10^{-4}$ is the leftover free electrons from after recombination.
Given the relatively small impact of the HeII reionization on $\tau_{\mathrm{reio}}$, the current uncertainties regarding its timeline, and the difficulty involved with its accurate modeling 56, 57, we opt to follow the conventional approach and include the second Helium reionization using the $\tanh$ model
$x_{\mathrm{e}}^{\mathrm{Tot}}=x_{\mathrm{e}}+\frac{Y_{\mathrm{He}}}{2C(1-Y_{\mathrm{He}})}\biggl{(}\tanh{\Bigl{(}\frac{z_{\mathrm{re}}^{\mathrm{HeII}}-z}{\Delta z^{\mathrm{HeII}}}\Bigr{)}}+1\biggr{)},$ (8)
where $z_{\mathrm{re}}^{\mathrm{HeII}}=3.5$ and $\Delta z^{\mathrm{HeII}}=0.5$ are the midpoint and duration of the second helium reionization, respectively. These choices are also the default values used by CLASS.
Shape universality and modeling
Figure 6: Universal shape of $x_{\mathrm{HI}}$. (Top) 128 simulated $x_{\mathrm{HI}}$ timelines (thin light gray lines) exhibit universality, after power-law transformations $\tilde{a}=(a/\alpha)^{\beta}$. The blue curves in all panels show our fitted analytic shape model, a composition of the Gompertz curve with a 5th-degree polynomial in Eqs. 2 and 3. (Middle) Time derivative of $x_{\mathrm{HI}}$, can be interpreted as a PDF if we view $x_{\mathrm{HI}}$ itself as the CDF. We first discovered the universality by translating and rescaling each $x_{\mathrm{HI}}$ using the mean and variance of its PDF, though now switch to the better approach that jointly fits the global shape and individual power-law parameters. We use the latter as target of symbolic regression. (Bottom) Timelines transformed by the inverse of Gompertz function, modeled in blue curve with a 5th-degree polynomial in Eq. 3.
We discovered the universality in the shape of $x_{\mathrm{HI}}$ timelines before attempting to build analytic model for it. Because $x_{\mathrm{HI}}$ varies monotonically between 1 and 0, we can view it as a CDF and derive its PDF, with which we can weigh the logarithmic scale factor $\ln a$ to compute its mean and standard deviations. It was immediately obvious to us that the $x_{\mathrm{HI}}$’s had a common shape to percent level, after translation by their means and rescaling by their standard deviations. However, given our broad parameter range in Simulations, some $x_{\mathrm{HI}}$’s have not reached 0 by the end of simulations, resulting in imperfect transformations hurting the universality.
To address this, we construct flexible models for the universal shape, and fit it jointly with individual transformation parameters of each $x_{\mathrm{HI}}$ timeline. We compose the Gompertz function $\mathrm{gomp}$, defined in Eq. 2, with a low-degree polynomial $P_{m}$, where $m=1,3,5,7$ progressively. We fit the composed shape to minimize the mean squared error (MSE) in 128 $x_{\mathrm{HI}}$’s and at 92 time points in each, and find the objective value improve with $m$ but only marginally from $P_{5}$ to $P_{7}$. Therefore, our final shape model is a composition of $\mathrm{gomp}$ and $P_{5}$, (see the lower panel of Figure 6), and has 6 parameters to fit.
As for the transformation parameters, as in the PDF approach, we use an affine transformation $\ln\tilde{a}=\beta(\ln a-\ln\alpha)$, or equivalently a power law in Eq. 4. Because each $x_{\mathrm{HI}}$ has its own parameters of $\alpha$ and $\beta$, we have in total $262=6+2\times 128$ parameters to determine in the joint fit. Figure 6 shows all 128 $x_{\mathrm{HI}}$ timelines and their universality after transformations. We can then use the fitted $\alpha$ and $\beta$ as the target for symbolic regression, to model their cosmological dependences on the other independent parameters.
Symbolic regression
Figure 7: Pareto front for symbolic regression (blue solid steps) illustrates the trade-off between regression accuracy and expression complexity. We further add the lower convex hull to aid model selection: each segment of the orange dotted line represents a power-law trade-off in this log-log plot, and every model that touches the convex hull is more economic – in the sense of accuracy gain at cost of complexity – than the nearby models above the segments. As an example, here we show the results of $\ln\alpha(\bm{\theta})$, where the complexity 11 point gives Eq. 5.
SR learns a model of data in the form of analytic expressions. Unlike traditional fittings that are restricted by their specific parameterization, SR searches in the vast function space of all expressions composed of specified operations, input variables, and free constants. It is NP-hard 58, 59 and typically need genetic or deep learning-based algorithms. In this work, we use the PySR package 60, 21 which performs SR optimizations using a multi-population genetic algorithm. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, it’d be great if you can refine/rewrite the SR and PySR intro
With PySR, we search for symbolic expressions that take the 6 parameters in Figure 5 as inputs and output $\ln\alpha$ or $\beta$, to minimize the MSE loss. Here, the search space is the set of expressions composed of 5 binary operators ($+$, --, $\cdot$, ///, and power function) and 2 unary operators ($\exp$ and $\ln$). Each expression naturally takes the form of a binary tree, and the total number of nodes is a measure of its complexity. We use 512 PySR populations each having 33 expressions, and optimize for 10000 iterations each with 10000 cycles. backgroundcolor=Bittersweet!20, linecolor=Bittersweet!50, inline]Miles, could you translate this into some brief description in the language of evolve-simplify-optimize loop, mutation, crossover, and tournament?
More complex expressions tend to fit more accurately, a trade-off typically visualized by the Pareto front, as shown in Figure 7. Better and more economic expressions can achieve a lower loss at moderate increase of complexity. To aid model selection, we use a heuristic that compares all expressions on the Pareto front globally, and only considers models that fare favorably in power-law trade-offs of the form
$\mathrm{loss}^{1-\gamma}\cdot\mathrm{complexity}^{\gamma}=\mathrm{const},\quad\forall\gamma\in(0,1).$ (9)
All such expressions lie on the lower convex hull of the Pareto front, as illustrated in Figure 7.
For the pivot $\ln\alpha$, we find complexity 11 is enough for the MSE loss to reach $10^{-4}$ (sub-percent level error on average), so there’s no need to use expressions more complex that that. However, for the tilt $\beta$, complexity 1, i.e. a constant, fit with a loss $\approx 0.1$ ($4\%$ error on average), while higher complexities can help but very slowly. Therefore, based on the economic heuristic, we choose expressions of complexity 11 and 1, for $\ln\alpha$ and $\beta$, respectively.
This preference for constant tilt could be a consequence of the astrophysical assumptions of reionization in 21cmFAST. A more complex simulation, such as one involving radiative transfer or different X-ray preheating 61, might reveal a different tilt.
[…]
Gompertz reionization Tanh reionization Parameter gomp 0 gomp 1 gomp 2 $\tanh$ 0 $\tanh$ 1 Planck PR31 $10^{9}A_{\mathrm{s}}$ 2.101(31) 2.088(12) 2.089(13) 2.099(30) 2.098(30) 2.100(30) $n_{\mathrm{s}}$ 0.9638(41) 0.9634(40) 0.9634(40) 0.9641(42) 0.9641(41) 0.9649(42) $\Omega_{\mathrm{c}}h^{2}$ 0.1201(12) 0.1203(11) 0.1202(11) 0.1201(12) 0.1201(12) 0.1200(12) $\Omega_{\mathrm{b}}h^{2}$ 0.02235(15) 0.02233(14) 0.02233(14) 0.02235(15) 0.02235(15) 0.02237(15) $\Omega_{\mathrm{m}}$ 0.3157(76) 0.3171(68) 0.3168(68) 0.3157(76) 0.3159(75) 0.3153(73) $\Omega_{\mathrm{m}}h^{2}$ 0.1430(11) 0.1433(10) 0.1432(10) 0.1430(11) 0.1431(11) 0.1430(11) $\Omega_{\Lambda}$ 0.6842(76) 0.6828(68) 0.6831(68) 0.6842(76) 0.6840(75) 0.6847(73) Age 13.800(23) 13.803(20) 13.803(22) 13.800(23) 13.800(23) 13.797(23) $H_{0}$ 67.32(55) 67.22(49) 67.24(50) 67.32(55) 67.31(54) 67.36(54) $100\theta_{\mathrm{X}}$ 1.04184(29) 1.04183(29) 1.04183(29) 1.04815(29) 1.04184(29) 1.04092(31) $\sigma_{8}$ 0.8109(59) 0.8092(45) 0.8092(46) 0.8108(60) 0.8105(59) 0.8111(60) $S_{8}$ 0.832(13) 0.832(13) 0.832(13) 0.832(13) 0.832(13) 0.832(13) $\tau_{\mathrm{reio}}$ 0.0546(76) 0.05115(61) 0.05140(63) 0.0543(75) 0.0538(74) 0.0544(73) $z_{\mathrm{re}}$ 6.81(68) 7.40 7.41 7.67(75) 7.62(74) 7.67(73)
Figure 8: Analysis of CMB data treating $\tau_{\mathrm{reio}}$ as a derived parameter using Eqs. 2, 3, 4 and 5 vs. sampling it with the conventional $\tanh$ model. Here gomp 1 combines our Gompertz universal shape with the rescaling pivot and tilt obtained via symbolic regression. Note that for gomp 1 we do not sample over $z_{\mathrm{re}}$ since Eq. 2 does not depend on it. The green (blue) contours correspond to the constraints obtained with our Gompertz universal shape ($\tanh$ 1 model).
Figure 9: Validation of our universal shape for $x_{\mathrm{HI}}$ in Eqs. 2 and 3 vs. the conventional $\tanh$ model using CMB data with sampling over $\tau_{\mathrm{reio}}$. Note that we have not eliminated $\tau_{\mathrm{reio}}$ from the analysis yet. Here, we use the standard choice of sampling over optical depth and obtain the corresponding reionization history via bisection. The green (blue) contours correspond to the constraints obtained with our gomp 0 ($\tanh$ 0) model.
Alternative mapping
The mapping between neutral hydrogen profiles and cosmology is not unique. This is because symbolic regression algorithms can be non-convergent. Moreover, the resulting symbolic expressions can depend on the data used for training, i.e. overfitting.
To investigate the impact of overfitting, we train an alternative mapping of reionization with cosmology using only half of the 21cmFAST $x_{\mathrm{HI}}$ profiles, which we refer as gomp 2. For the tilt, once again we find a constant $\beta=8.331$, remarkably close to the value obtained using the full simulated data (8.290). And for the pivot, we obtain
$\ln\alpha(\bm{\theta})=(1.039+\sigma_{8})(\Omega_{\mathrm{b}}-\Omega_{\mathrm{m}}h-n_{\mathrm{s}}),$ (10)
with a training loss of $1.2\times 10^{-4}$. Now we can test this on the other half of our $x_{\mathrm{HI}}$ sample and get a validation loss of $1.4\times 10^{-4}$, only slightly larger than the training loss implying we are safe from overfitting.
Unsurprisingly, Eq. 10 recovers the cosmological trends present in Eq. 5. Specifically, increasing $n_{\mathrm{s}}$, $\Omega_{\mathrm{m}}$, and $\sigma_{8}$ leads to an earlier reionization scenario, while a higher value of $\Omega_{\mathrm{b}}$ slightly delays reionization. Furthermore, the impact of the Hubble constant is once again linked to the matter density, albeit not in the same manner as in Eq. 5. This is likely evidence that the role of $h$ is limited, and PySR considers its influence in conjunction with $\Omega_{\mathrm{m}}$ without increasing the complexity of the analytic expression.
Replacing Eq. 5 with Eq. 10, we rerun the analysis following the same approach, and summarize the results in Table 1. Comparing these findings to our previous results, we observe strikingly similar values for all cosmological parameters, albeit with smaller deviations from the Planck results. The consistency between our results using the full $x_{\mathrm{HI}}$ sample and those using only half of it indicates that our symbolic expression mappings perform robustly and are not biasing the parameter inferences.
MCMC inference
Table 1 summarizes the results obtained by performing MCMC Bayesian inference with Cobaya for the models considered throughout this work. We also include the Planck results 1 for reference. In total, we run three Gompertz reionization and two $\tanh$ reionization models, where $\tanh$ 0, Planck PR3, and gomp 0 sample over the typical 6 cosmological parameters (see shaded parameters in Table 1). In contrast, gomp 1, and gomp 2 sample over $\bm{\theta}=\{\sigma_{8},n_{\mathrm{s}},\Omega_{\mathrm{b}}h^{2},\Omega_{\mathrm{c}}h^{2},h\}$. Similarly, $\tanh$ 1 samples over the same parameters as gomp 1 but with the inclusion of $z_{\mathrm{re}}$, and agrees with $\tanh$ 0 as the expected result from this sanity check.
There is an apparent discrepancy between the different models in $\theta_{\mathrm{x}}$, a proxy for the angular scale of the acoustic oscillations ($\theta_{*}$). The difference is due to our use of $100\theta_{\mathrm{s}}$, which corresponds to the peak scale parameter defined exactly as the ratio of the sound horizon divided by the angular diameter at decoupling, with decoupling time given by the maximum of the visibility function, i.e. the standard choice for CLASS. In contrast, the Planck collaboration reports $100\theta_{\mathrm{MC}}$, which is given by Eq. (6) in 62, and corresponds to the standard choice in CosmoMC26.
Although Table 1 does not display an error in the midpoint of reionization in the gomp 1 and gomp 2 models, it is important to note that an error does exist. Unfortunately, it cannot be automatically computed with Cobaya because gomp 1 and gomp 2 do not sample over reionization parameters. Therefore, $z_{\mathrm{re}}$ has no meaning in our Gompertz CLASS [^3] This oversight can be corrected; however, it would require re-running the chains. for models that take advantage of Eq. 1.
Remarkably, the inferred value of $S_{8}$ remains consistent across all reionization models, even when $\sigma_{8}$ and $\Omega_{\mathrm{m}}$ vary between models. This trend suggests that different reionization models neither alleviate nor exacerbate the $S_{8}$ tension. While this observation holds for the Planck PR3 data, we anticipate that incorporating low-redshift Baryon Acoustic Oscillation (BAO) data may affect this conclusion. Future work will explore the implications of Gompertz reionization for the joint analysis of CMB and low-redshift data.
Supplementary information
No.
Acknowledgements
This work is supported by The Major Key project of PCL. The authors acknowledge PCL’s Cloud Brain for providing computational and data storage resources that have contributed to the results reported within this paper.
Declarations
Authors contributions
P.M.C. and Y.L. contributed equally to all stages of the project. M.C. contributed to the implementation of the symbolic regression algorithm.
Correspondence and requests for materials
Should be addressed to Paulo Montero-Camacho and Yin Li.
Code availability
[…]
Funding
Not applicable.
Competing interests
The authors declare no competing interests.
Ethics approval
Not applicable.
Consent to participate
Not applicable.
Consent for publication
Not applicable.